OpenStudy (anonymous):

Help please. The length of the shadow of a building is 100 meters as shown below. What is the height of the building? 50 m 100 m 200 m 55 m

OpenStudy (anonymous):

i think that the way you will figure this problem out, is by using trigonometric functions.

OpenStudy (anonymous):

can you please explain to me how to do this please?

OpenStudy (anonymous):

umm ok im not 100% sure but ill try let me draw this.

OpenStudy (anonymous):

okay thanks

OpenStudy (anonymous):



OpenStudy (anonymous):

so i think u will either have to use sin, cos or tan. let me check to see

OpenStudy (anonymous):

wait......there is no angle.

OpenStudy (anonymous):

ok there is a 45 degree angle i see.

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

ok you will use the cos (cosine) function

OpenStudy (anonymous):

f(x) = a*cos(bx + c) + d that?

OpenStudy (anonymous):

\[\cos 45 = \frac{ 100 }{ x }\]

OpenStudy (anonymous):

How do i solve it?

OpenStudy (anonymous):

wait i might be wrong.

OpenStudy (anonymous):

answer choices 50 m 100 m 200 m 55 m I think its B

OpenStudy (anonymous):

i mean D

OpenStudy (anonymous):

it may be 200m

OpenStudy (anonymous):

but wait let me check

OpenStudy (anonymous):

how did you get 200

OpenStudy (anonymous):

i didnt lol wait i have to try to figure it out. this problem is still unsolved at this second.

OpenStudy (anonymous):

lol okay im trying it too

OpenStudy (anonymous):

it is 100

OpenStudy (anonymous):

one sec

OpenStudy (anonymous):

how did you get 100 an you please show the steps you did?

OpenStudy (anonymous):

give me 5 minutes and i will get you an explanation.

OpenStudy (anonymous):

okay thank you.

OpenStudy (anonymous):

i dont know. someone else will clarify. just wait here ill try to find someone.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

are you talking calculus 1?

OpenStudy (anonymous):

no Geometry

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

im checking my text book

OpenStudy (anonymous):

okau if i go offline its not my fault my laptop has a virus and it tends to turn on and off by itself

OpenStudy (anonymous):

ok i think i may have some clarification

OpenStudy (anonymous):

okay lol

OpenStudy (anonymous):

ok i think the correct method is the tan method. I will draw this out and lets see what we get.

OpenStudy (anonymous):